OpenStudy (anonymous):

Write the equation of the line that passes through (-3, 4) with a slope of 2 in point-slope form. Answers: 1) y - 4 = 2(x + 3) 2) y + 4 = 2(x - 3) 3) y + 3 = 2(x - 4) 4) y - 3 = 2(x + 4) can anyone help?

OpenStudy (anonymous):

?

OpenStudy (bahrom7893):

wait never mind that

OpenStudy (bahrom7893):

y-4=2(x+3)

OpenStudy (bahrom7893):

y-y1=M(x-x1)

OpenStudy (saifoo.khan):

1) y - 4 = 2(x + 3)

OpenStudy (anonymous):

so number one is the correct answer?

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

y=2x+10

OpenStudy (saifoo.khan):

\[\Huge Yea!\]

OpenStudy (anonymous):

okay thanks ill fan you !